Broadening Your Perspective 5-2 (Part Level Submission) The condensed income statement for the Peri and Paul partnership for 2017 is as follows PERI AND PAUL COMPANY Income Statement For the Year Ended December 31, 2017 Sales (240,000 units) Cost of goods sold Gross profit $1,200,000 900,000 400,000 $280,000 150,000 Selling Administrative 430,000 Net loss $(30,000) A cost behavior analysis indicates that 60% of the cost of goods sold are variable 45% of the selling expenses are variable, and 37% of the administrative expenses are variable (Round to nearest unit, dollar, and percentage, where necessary. Use the CVP income statement format in computing profits.) Compute the break-even point in total sales dollars and in units for 2017. (Round intermediate calculations to 2 decimal places, e.g. 0.25 and final answers to 0 decimal places, e.g. 2,520.) Break-even point in dollars Break-even point in units units Attempts: 0 of 2 used SAVEFOR LATER SUBMIT ANSWER
Answer #1
Apportionment of cost
Variable Fixed
Cost of goods sold 480000 320000
Seling expense 126000 154000
Admin Expense 55500 94500
Total 661500 568500
Sales 1200000
Total Variable Cost 661500
Contribution 538500
Number of units 240000
Contribution per unit 2.24375
CM ratio: Contribution/ Sales *100
538500/1200000*100 = 44.875%
Break even in $: Fixed cost / CM ratio
$ 568500 /44.875% = $1266,852
Break even in Units: Fixed cost / CM per unit
$ 568500 /2.24375 = 253,371 units